OpenStudy (anonymous):

Help? 2t+ u = v for t

OpenStudy (anonymous):

2t+u=v 2t=v/u --->2v/u i think

OpenStudy (anonymous):

Isolate the t. 2t+u=v Subtract u from both sides. 2t=v-u Divide 2 on both sides t=v-u/2